OpenStudy (stormoshawty):

What is the x-value of the solution? -x + 2y = 6 6y = x + 18

OpenStudy (anonymous):

\(-x+2y=6\) can be rewritten as \(x=2y-6\)

OpenStudy (anonymous):

take the \(x\) in the second equation and replace it by \(2y-6\) to get \[6y=2y-6+18\] or \[6y=2y+12\] then you can solve for \(y\) in two steps